The random variable X is a payout table of a casino slot machine. The probability mass function is given:

X -5 0 2 10 20 40 60 1000
Probability 0.65 0.159 0.10 0.05 0.02 0.01 0.01 0.001

Find the variance for the payout .
Answer #1

Mean = X * P(X)

= -5 * 0.65 + 0 * 0.159 + 2 * 0.10 + 10 * 0.05 + 20 * 0.02 + 40 * 0.01 + 60 * 0.01 + 1000 * 0.001

= -0.15

Variance = X2 * P(X) - Mean2

= [ -52 * 0.65 + 02 * 0.159 + 22 * 0.10 + 102 * 0.05 + 202 * 0.02 + 402 * 0.01 + 602 * 0.01 + 10002 * 0.001 - (-0.15)2

= 1081.6275